OpenStudy (anonymous):

Box A has a volume of 48 cubic meters. Box B is similar to box A. To create box B, box A's dimensions were doubled. What is the volume of box B? (1 point) Select one: a. 24 m3 b. 96 m3 c. 288 m3 d. 384 m3

OpenStudy (anonymous):

@Nurali

OpenStudy (whpalmer4):

The volume of a rectangular prism with dimensions \(l,w,h\) is \(V = l*w*h\). What happens to \(V\) if you replace \(l\) with \(2l\), \(h\) with \(2h\), and \(w\) with \(2w\)?

OpenStudy (anonymous):

V = 6 right?

OpenStudy (whpalmer4):

No, \[V = (2l)*(2h)*(2w) = 2*2*2*l*h*w =\]

OpenStudy (whpalmer4):

The problem author did put in a corresponding answer to trap anyone who made that mistake, however :-)

OpenStudy (whpalmer4):

the new volume would be \[8lhw\] compared to the old \[lhw\]right? if \[48 \text{ m}^3 = l*h*w\]then what is \[8*l*h*w=\]?

OpenStudy (anonymous):

so do i multiply something

OpenStudy (whpalmer4):

Yes. the new volume is the same as the old volume * 8, right? \[\text{old volume} = 48 \text{ m}^3 = l*h*w\] \[\text {new volume} = 8*l*h*w = 8 * (48\text{ m}^3) =\]

OpenStudy (anonymous):

so what do you get

OpenStudy (whpalmer4):

the new volume is 8* the old volume. you should be able to work that out...